OpenStudy (chaise):

Solve: cos(x/4)=1/sqrt(2) It's been a while since I have done this stuff, can anyone help me, or atleast give me a hint? Thanks in advance.

OpenStudy (chaise):

\[\cos(\frac{x}{4})=\frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

At what angle cos becomes 1/root(2) ??

OpenStudy (chaise):

45

OpenStudy (anonymous):

Yes so: \[\cos(\frac{ x}{4}) = \cos(\frac{\pi}{4})\] So what can be x here???

OpenStudy (chaise):

Oh Cool, that means x must be pi, that helps a lot, thanks ^_^

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

Are you looking for the general solution of this??

OpenStudy (chaise):

No thanks, just the solutions between -2pi>x>2pi, but there is only two solutions. +/- pi :) Thanks, I may have another shortly :)

OpenStudy (anonymous):

It is true that one value of x is \( \pi\) but there can be more values of x..

OpenStudy (anonymous):

For knowledge I will tell you more that : For \(cos \theta = cos \alpha\) the general solution is given by: \(\theta = 2n \pi + \alpha\), where \(n\) is an integer... So for this question: \[\cos(\frac{x}{4}) = \cos(\frac{\pi}{4}) \implies \frac{x}{4} = 2n \pi + \frac{\pi}{4}\] \[\large \implies x = 8n \pi + \pi\] Put n = 0 here you will get: \(x = \pi\) Similarly put n = 1, 2 or -1, -2 you will get infinite values for x...

OpenStudy (chaise):

I know, I'm only looking for the solutions between -7>x>7 on the x axis. Thanks anyway :)

OpenStudy (anonymous):

Welcome dear..